# Combinatorics of Essential Sets for Positroids

### v2

@@ 4 Combinatorial characterization of essential sets
The following result is a generalization of Theorem 4.1 from [5] to the setting of bounded affine permutations and our definition of the rank of an essential set. We retain the same name, which stands for CHaracterization of ESsential Sets. The use of cyclic intervals and bounded affine permutations significantly alters the third condition in the theorem. The rank interpretation of essential sets allows for an elegant proof of the result. We divide its proof into Lemmas 4.1 and 4.2. Lemma 4.1 demonstrates that we can use a family $\mathcal{E}^{r}$ satisfying properties (E1), (E2), and (E3) to construct a function on cyclic intervals which shares many properties with the rank function of a positroid.
@@
Let $\mathcal{E}^{r}$ be a set of cyclic intervals labeled by a non-negative integer satisfying the properties (E1), (E2), and (E3) from Theorem 3.13. Let $\operatorname{CI}_{n}$ denote the family of cyclic intervals of $[n]$ and $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined by:
$r([i,j])=\min\{|[i,j]|,r+|[i,j]\setminus I|\mid(r,I)\in\mathcal{E}^{r}\}.$
@@
    $r(I)=r$ for every $(r,I)\in\mathcal{E}^{r}$ and the minimum is uniquely achieved by $(r,I)$;
@@
    $[i,j]\in\operatorname{CI}_{n}$ is such that $r([i,j])=r([i+1,j])=r([i,j-1])=r([i-1,j])-1=r([i,j+1])-1$ if and only if $(r([i,j]),[i,j])\in\mathcal{E}^{r}$.
@@
(r1): Let $(r,I)\in\mathcal{E}^{r}$. By property (E1), we have $r<|I|$. Suppose by contradiction that there exists $(r_{2},I_{2})\in\mathcal{E}^{r}$ such that $r_{2}+|I_{2}\setminus I|\leq r$. By (E2), $I_{2}$ neither contains nor is contained by $I$. Hence, let $(r_{3},I_{3})$ be a minimal set with the smallest label in $\mathcal{E}^{r}$ containing $I\cup I_{2}$, and let $(r_{4},I_{4})$ be a maximal set with the largest label contained in $I\cap I_{2}$. Suppose $I$ and $I_{2}$ intersect in a cyclic interval. By (E3), $r+r_{2}\geq r_{3}+r_{4}+|(I\cap I_{2})\setminus I_{4}|$. Moreover, $r_{4}+|(I\cap I_{2})\setminus I_{4}|+|I\setminus I_{2}|=r_{4}+|I\setminus I_{4}|\geq r_{2}+|I\setminus I_{2}|$, where the last equality holds by assuming that the minimum of the function $r$ is achieved by $(r_{2},I_{2})$. Then we can write $r+r_{2}+|I\setminus I_{2}|\geq r_{3}+r_{2}+|I\setminus I_{2}|$, hence $r\geq r_{3}$. Since $I\subsetneq I_{3}$, by (E2), this is not possible. Finally, suppose $I$ and $I_{2}$ intersect in two cyclic intervals. Then $r_{2}+|I\setminus I_{2}|=r_{2}+|[1,n]\setminus I_{2}|\geq r_{2}+k-r_{2}+k>r$, where we used (E2). It follows that for every $(r_{2},I_{2})\in\mathcal{E}^{r}\setminus{(r,I)}$, $r_{2}+|I\setminus I_{2}|>r$, hence (r1) holds.
(r2): Property (r2) holds by definition of the function $r$ and condition (E1) on the family $\mathcal{E}^{r}$.
@@
  • Suppose $r([i,j])=|[i,j]|$ and $r([i-1,j])=r+|[i-1,j]\setminus I|$ for some $(r,I)\in\mathcal{E}^{r}$. Then $i-1\in I$ and $r+|[i,j]\setminus I|\geq|[i,j]|$, hence $r\geq|I\cap[i,j]|$. It follows that $r([i-1,j])=r+|[i-1,j]\setminus I|\geq|I\cap[i,j]|+|[i-1,j]\setminus I|=|[i,j]|=r([i,j])$. Moreover $r([i-1,j])\leq|[i-1,j]|=r([i,j])+1$.
@@
(r4): Suppose $[i,j]\subsetneq[n]$ is such that $r([i,j])=r([i+1,j])=r([i,j-1])=r([i-1,j])-1=r([i,j+1])-1$ holds. Let $(r,I)\in\mathcal{E}^{r}$ such that $r([i,j])=r+|[i,j]\setminus I|$. Note that by assumption, $i-1,j+1\not\in I$, hence $I\subseteq[i,j]$. Suppose $I\neq[i,j]$ and assume that $i\not\in I$. Then $r([i+1,j])\leq r+|[i+1,j]\setminus I|<r+|[i,j]\setminus I|=r([i,j])$, a contradiction. In the same way we need to have that $j\in I$. Since $I$ is a cyclic interval, it follows that $I=[i,j]$. By (r1), $r([i,j])=r$, hence $(r([i,j]),[i,j])\in\mathcal{E}^{r}$.
Suppose now $(r,[i,j])\in\mathcal{E}^{r}$. By (r1), $r([i,j])=r$ and the minimum is uniquely achieved by $(r,I)$. Hence, for every $(r_{2},I_{2})\in\mathcal{E}^{r}$, $r_{2}+|[i,j]\setminus I_{2}|\geq r+1$ and $r([i+1,j])\geq r$, $r([i,j-1])\geq r$. Since they are both contained in $[i,j]$, it follows that $r([i,j])=r([i+1,j])=r([i,j-1])=r$. Moreover, $r_{2}+|[i-1,j]\setminus I_{2}|\geq r+1$ and $r_{2}+|[i,j+1]\setminus I_{2}|\geq r+1$, and $r+|[i-1,j]\setminus[i,j]|=r+1=r+|[i,j+1]\setminus[i,j]|$, hence we can conclude that (r4) holds. ∎
@@
Let $\mathcal{E}^{r}$ be a family of labeled cyclic intervals satisfying (E1), (E2), and (E3) from Theorem 3.13. Let $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined in Lemma 4.1 and let $\pi:\mathbb{Z}\to\mathbb{Z}$ be defined as
@@
Suppose $i,j\in\mathbb{Z}$, with $i<j$, are such that $\pi(i)=\pi(j)=\ell$. Then $r([j,\ell])=r([j+1,\ell])>r([j+1,\ell-1])$. By property (r3) of the function $r$, $r([j+1,\ell-1])=r([j,\ell])-1$. We have that $r([j,\ell])\geq r([j,\ell-1])\geq r([j+1,\ell-1])=r([j,\ell])-1$. Suppose the first inequality is strict, then $r([j,\ell-1])=r([j+1,\ell-1])$, contradicting the minimality assumption on $\ell$. Hence $r([j,\ell])=r([j,\ell-1])$. In the same way, $r([i,\ell])=r([i,\ell-1])$. Hence there are $(r_{i},I),(r_{j},J)\in\mathcal{E}^{r}$ such that $[i,\ell]\subseteq I$, $[j,\ell]\subseteq J$ and $r_{i}=r([i,\ell])$, $r_{j}=r([j,\ell])$. By definition of $\pi(i)=\ell$, it follows that $r([i+1,\ell-1])=r_{i}-1$. Suppose $r([i+1,\ell-1])=r_{2}+|[i+1,\ell-1]\setminus I_{2}|$ for some $(r_{2},I_{2})\in\mathcal{E}^{r}$, then $I_{2}$ is contained in $[i+1,\ell-1]$ since if I add $i$ or $\ell$, the value of $r$ strictly increases. Then, $r_{2}+|[i+1,\ell-1]\setminus I_{2}|\geq r_{2}+|I\setminus I_{2}|\geq r$ by property (E2), contradicting the assumption. Finally, note that $|[i+1,\ell-1]|\geq|[i+1,\ell-1]|+r^{\prime}-|[j,\ell-1]|\geq r^{\prime}+|[i+1,\ell-1]\setminus J|=r^{\prime}+|[i+1,\ell]\setminus J|\geq r$. Hence, we get a contradiction, and the function $\pi$ is injective. Moreover, since $\pi$ is bounded and periodic, it follows that $\pi$ is a bijection. ∎
@@
Let $\mathcal{E}^{r}$ be a set of cyclic intervals labeled by a non-negative integer. Then $\mathcal{E}^{r}$ is the ranked essential family of a positroid if and only if $\mathcal{E}^{r}$ satisfies the properties (E1), (E2), and (E3) from Theorem 3.13.
@@
If $\mathcal{E}^{r}$ is the ranked essential family of a positroid, then (E1), (E2), and (E3) hold as shown in Theorem 3.13.
Suppose now $\mathcal{E}^{r}$ is family of labeled cyclic intervals satisfying (E1), (E2), and (E3). Let $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined in Lemma 4.1 and let $\pi:\mathbb{Z}\to\mathbb{Z}$ be the bounded affine permutation constructed in Lemma 4.2. We want to prove that the positroid $\mathcal{P}$ associated to $\pi$ is such that rk ([$\textrm{rk }([i,j])=r([i,j])$ for each cyclic interval $[i,j]$. Then, by Theorem 3.8 and property (r4) from Lemma 4.1, it follows that $\mathcal{E}^{r}$ is the ranked essential family of the positroid $\mathcal{P}$. We will prove by induction on the size of the cyclic interval that the rank ‘rk ’ and ‘$r$’ agree on each cyclic interval. For $i\in[n]$, we have
@@
Case 1. If $r([i,j])<\textrm{rk }([i,j])$, since both $r$ and rk are satisfying the property (r3) from Lemma 4.1, $r([i,j])=r([i+1,j])$, hence $\pi(i)\leq j$. Moreover, rk ([$\textrm{rk }([i,j])=\textrm{rk }([i+1,j])+1$, i.e., $i$ is independent of $i=1,\dots,j$, hence $\pi(i)>j$ and we get a contradiction.
@@
Case 2.1. Suppose $\pi(j)=j$, then $r([j,j])=0$ hence there is $(0,I)\in\mathcal{E}^{r}$ such that $j\in I$. Suppose $r([i,j-1])=r^{\prime}+|[i,j-1]\setminus I^{\prime}|$ for some $(r^{\prime},I^{\prime})\in\mathcal{E}^{r}\cup\{(0,\emptyset)\}$. First assume that $I\cap I^{\prime}\neq\emptyset$. Let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{r}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{r}$ contained in $I\cap I^{\prime}$. Then by (E3.2) $0+r^{\prime}\geq r_{3}+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|$. Then $r^{\prime}+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$. Hence $r([i,j-1])=r_{3}+|[i,j-1]\setminus I_{3}|=r([i,j])$, where the last equality holds since $j\in I_{3}$. Assume now that $I\cap I^{\prime}\neq\emptyset$. Similarly, let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{r}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{r}$ contained in the minimal cyclic interval containing $I\cup I^{\prime}$ in $[i,j]$ $C$, minus $I\cup I^{\prime}$. Then by (E3.1), $0+r^{\prime}\geq r_{3}-r_{4}-|C\setminus I_{4}|$. Then $r^{\prime}+|[i,j-1]\setminus I^{\prime}|\geq r_{3}-r_{4}-|C\setminus I_{4}|+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$ and as before we get that $r([i,j])=r([i,j-1])$ and we get that $r$ and $\operatorname{rk}$ agree on $[i,j]$.
Case 2.2. Suppose $\pi(j)>j$. Since rk ([$\textrm{rk }([i,j-1])=\textrm{rk }([i,j])$, by Proposition 2.7, there is some $h\in[i+1,j-1]$ such that $\pi(h)=j$, i.e., $r([h,j])=r([h+1,j])=r([h,j-1])$. By (r4), there exists some $(r,I)$ such that $[h,j]\subseteq I$ and $r=r([h,j])=\textrm{rk }([h,j])$. Let $(r^{\prime},I^{\prime})\in\mathcal{E}^{r}\cup\{(0,\emptyset)\}$ be such that $r([i,j-1])=r^{\prime}+|[i,j-1]\setminus I^{\prime}|$. Suppose $I\cap I^{\prime}\neq\emptyset$, the other case will be equivalent. Let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{r}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{r}$ contained in $I\cap I^{\prime}$. Then by (E3), $r^{\prime}+|[i,j-1]\setminus I^{\prime}|=r^{\prime}+|[i,j-1]\setminus(I\cup I^{\prime})|+|I\setminus I^{\prime}|\geq r^{\prime}+|[i,j-1]\setminus I_{3}|+|I\setminus I^{\prime}|\geq r+r^{\prime}-|I\cap I^{\prime}|+|[i,j-1]\setminus I_{3}|\geq r_{3}+|[i,j-1]\setminus I_{3}|+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|-|I\cap I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$. Hence $r([i,j-1])=r_{3}+|[i,j-1]\setminus I_{3}|=r_{3}+|[i,j]\setminus I_{3}|=r([i,j])$, since $j\in I_{3}$. ∎
@@
If $\mathcal{E}^{r}$ is a family of labeled cyclic interval satisfying (E1), (E2), and (E3) we can recover the bounded affine permutation of the positroid defined by $\mathcal{E}^{r}$ by applying Algorithm 1 to $E^{r}=\{(r,(i,j))\mid(r,[i,j+i-1])\in\mathcal{E}^{r}\}$.

### v3

@@ 4 Combinatorial characterization of essential sets
The following result is a generalization of Theorem 4.1 from [6] to the setting of bounded affine permutations and our definition of the rank of an essential set. We retain the name “chess theorem”, which stands for CHaracterization of ESsential Sets. The use of cyclic intervals and bounded affine permutations significantly alters the third condition in the theorem. The rank interpretation of essential sets allows for an elegant proof of the result. We divide its proof into Lemmas 4.1 and 4.2. Lemma 4.1 demonstrates that we can use a family $\mathcal{E}^{\mathrm{r}}$ satisfying properties (E1), (E2), and (E3) to construct a function on cyclic intervals which shares many properties with the rank function of a positroid.
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a set of cyclic intervals labeled by a non-negative integer satisfying the properties (E1), (E2), and (E3) from Theorem 3.13. Let $\operatorname{CI}_{n}$ denote the family of cyclic intervals of $[n]$ and $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined by:
$r([i,j])=\min\{|[i,j]|,r+|[i,j]\setminus I|\mid(r,I)\in\mathcal{E}^{\mathrm{r}}\}.$
@@
    $r(I)=r$ for every $(r,I)\in\mathcal{E}^{\mathrm{r}}$ and the minimum is uniquely achieved by $(r,I)$;
@@
    $[i,j]\in\operatorname{CI}_{n}$ is such that $r([i,j])=r([i+1,j])=r([i,j-1])=r([i-1,j])-1=r([i,j+1])-1$ if and only if $(r([i,j]),[i,j])\in\mathcal{E}^{\mathrm{r}}$.
@@
(r1): Let $(r,I)\in\mathcal{E}^{\mathrm{r}}$. By property (E1), we have $r<|I|$. Suppose for contradiction that there exists $(r_{2},I_{2})\in\mathcal{E}^{\mathrm{r}}$ such that $r_{2}+|I_{2}\setminus I|\leq r$. By (E2), $I_{2}$ neither contains nor is contained by $I$. Hence, let $(r_{3},I_{3})$ be a minimal set with the smallest label in $\mathcal{E}^{\mathrm{r}}$ containing $I\cup I_{2}$, and let $(r_{4},I_{4})$ be a maximal set with the largest label contained in $I\cap I_{2}$. Suppose $I$ and $I_{2}$ intersect in a cyclic interval. By (E3), $r+r_{2}\geq r_{3}+r_{4}+|(I\cap I_{2})\setminus I_{4}|$. Moreover, $r_{4}+|(I\cap I_{2})\setminus I_{4}|+|I\setminus I_{2}|=r_{4}+|I\setminus I_{4}|\geq r_{2}+|I\setminus I_{2}|$, where the last equality holds by assuming that the minimum of the function $r$ is achieved by $(r_{2},I_{2})$. Then we can write $r+r_{2}+|I\setminus I_{2}|\geq r_{3}+r_{2}+|I\setminus I_{2}|$, hence $r\geq r_{3}$. Since $I\subsetneq I_{3}$, by (E2), this is not possible. Finally, suppose $I$ and $I_{2}$ intersect in two cyclic intervals. Then $r_{2}+|I\setminus I_{2}|=r_{2}+|[1,n]\setminus I_{2}|\geq r_{2}+k-r_{2}+k>r$, where we used (E2). It follows that for every $(r_{2},I_{2})\in\mathcal{E}^{\mathrm{r}}\setminus{(r,I)}$, $r_{2}+|I\setminus I_{2}|>r$, hence (r1) holds.
(r2): Property (r2) holds by definition of the function $r$ and condition (E1) on the family $\mathcal{E}^{\mathrm{r}}$.
@@
  • Suppose $r([i,j])=|[i,j]|$ and $r([i-1,j])=r+|[i-1,j]\setminus I|$ for some $(r,I)\in\mathcal{E}^{\mathrm{r}}$. Then $i-1\in I$ and $r+|[i,j]\setminus I|\geq|[i,j]|$, hence $r\geq|I\cap[i,j]|$. It follows that $r([i-1,j])=r+|[i-1,j]\setminus I|\geq|I\cap[i,j]|+|[i-1,j]\setminus I|=|[i,j]|=r([i,j])$. Moreover $r([i-1,j])\leq|[i-1,j]|=r([i,j])+1$.
@@
(r4): Suppose $[i,j]\subsetneq[n]$ is such that $r([i,j])=r([i+1,j])=r([i,j-1])=r([i-1,j])-1=r([i,j+1])-1$ holds. Let $(r,I)\in\mathcal{E}^{\mathrm{r}}$ such that $r([i,j])=r+|[i,j]\setminus I|$. Note that by assumption, $i-1,j+1\not\in I$, hence $I\subseteq[i,j]$. Suppose $I\neq[i,j]$ and assume that $i\not\in I$. Then $r([i+1,j])\leq r+|[i+1,j]\setminus I|<r+|[i,j]\setminus I|=r([i,j])$, a contradiction. In the same way we need to have that $j\in I$. Since $I$ is a cyclic interval, it follows that $I=[i,j]$. By (r1), $r([i,j])=r$, hence $(r([i,j]),[i,j])\in\mathcal{E}^{\mathrm{r}}$.
Suppose now $(r,[i,j])\in\mathcal{E}^{\mathrm{r}}$. By (r1), $r([i,j])=r$ and the minimum is uniquely achieved by $(r,I)$. Hence, for every $(r_{2},I_{2})\in\mathcal{E}^{\mathrm{r}}$, $r_{2}+|[i,j]\setminus I_{2}|\geq r+1$ and $r([i+1,j])\geq r$, $r([i,j-1])\geq r$. Since they are both contained in $[i,j]$, it follows that $r([i,j])=r([i+1,j])=r([i,j-1])=r$. Moreover, $r_{2}+|[i-1,j]\setminus I_{2}|\geq r+1$ and $r_{2}+|[i,j+1]\setminus I_{2}|\geq r+1$, and $r+|[i-1,j]\setminus[i,j]|=r+1=r+|[i,j+1]\setminus[i,j]|$, hence we can conclude that (r4) holds. ∎
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a family of labeled cyclic intervals satisfying (E1), (E2), and (E3) from Theorem 3.13. Let $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined in Lemma 4.1 and let $\pi:\mathbb{Z}\to\mathbb{Z}$ be defined as
@@
Suppose $i,j\in\mathbb{Z}$, with $i<j$, are such that $\pi(i)=\pi(j)=\ell$. Then $r([j,\ell])=r([j+1,\ell])>r([j+1,\ell-1])$. By property (r3) of the function $r$, $r([j+1,\ell-1])=r([j,\ell])-1$. We have that $r([j,\ell])\geq r([j,\ell-1])\geq r([j+1,\ell-1])=r([j,\ell])-1$. Suppose the first inequality is strict, then $r([j,\ell-1])=r([j+1,\ell-1])$, contradicting the minimality assumption on $\ell$. Hence $r([j,\ell])=r([j,\ell-1])$. In the same way, $r([i,\ell])=r([i,\ell-1])$. Hence there are $(r_{i},I),(r_{j},J)\in\mathcal{E}^{\mathrm{r}}$ such that $[i,\ell]\subseteq I$, $[j,\ell]\subseteq J$ and $r_{i}=r([i,\ell])$, $r_{j}=r([j,\ell])$. By definition of $\pi(i)=\ell$, it follows that $r([i+1,\ell-1])=r_{i}-1$. Suppose $r([i+1,\ell-1])=r_{2}+|[i+1,\ell-1]\setminus I_{2}|$ for some $(r_{2},I_{2})\in\mathcal{E}^{\mathrm{r}}$, then $I_{2}$ is contained in $[i+1,\ell-1]$ since if I add $i$ or $\ell$, the value of $r$ strictly increases. Then, $r_{2}+|[i+1,\ell-1]\setminus I_{2}|\geq r_{2}+|I\setminus I_{2}|\geq r$ by property (E2), contradicting the assumption. Finally, note that $|[i+1,\ell-1]|\geq|[i+1,\ell-1]|+r^{\prime}-|[j,\ell-1]|\geq r^{\prime}+|[i+1,\ell-1]\setminus J|=r^{\prime}+|[i+1,\ell]\setminus J|\geq r$. Hence, we get a contradiction, and the function $\pi$ is injective. Moreover, since $\pi$ is bounded and periodic, it follows that $\pi$ is a bijection. ∎
@@
Let $\mathcal{E}^{\mathrm{r}}$ be a set of cyclic intervals labeled by a non-negative integer. Then $\mathcal{E}^{\mathrm{r}}$ is the ranked essential family of a positroid if and only if $\mathcal{E}^{\mathrm{r}}$ satisfies the properties (E1), (E2), and (E3) from Theorem 3.13.
@@
If $\mathcal{E}^{\mathrm{r}}$ is the ranked essential family of a positroid, then (E1), (E2), and (E3) hold as shown in Theorem 3.13.
Suppose now $\mathcal{E}^{\mathrm{r}}$ is family of labeled cyclic intervals satisfying (E1), (E2), and (E3). Let $r:\operatorname{CI}_{n}\to\mathbb{Z}$ be the function defined in Lemma 4.1 and let $\pi:\mathbb{Z}\to\mathbb{Z}$ be the bounded affine permutation constructed in Lemma 4.2. We want to prove that the positroid $\mathcal{P}$ associated to $\pi$ is such that rk ([$\textrm{rk }([i,j])=r([i,j])$ for each cyclic interval $[i,j]$. Then, by Theorem 3.8 and property (r4) from Lemma 4.1, it follows that $\mathcal{E}^{\mathrm{r}}$ is the ranked essential family of the positroid $\mathcal{P}$. We will prove by induction on the size of the cyclic interval that the rank ‘rk ’ and ‘$r$’ agree on each cyclic interval. For $i\in[n]$, we have
@@
Case 1. If $r([i,j])<\textrm{rk }([i,j])$, since both $r$ and rk are satisfying the property (r3) from Lemma 4.1, $r([i,j])=r([i+1,j])$, hence $\pi(i)\leq j$. Moreover, rk ([$\textrm{rk }([i,j])=\textrm{rk }([i+1,j])+1$, i.e. $i$ is independent of $i=1,\dots,j$, hence $\pi(i)>j$ and we get a contradiction.
@@
Case 2.1. Suppose $\pi(j)=j$, then $r([j,j])=0$ hence there is $(0,I)\in\mathcal{E}^{\mathrm{r}}$ such that $j\in I$. Suppose $r([i,j-1])=r^{\prime}+|[i,j-1]\setminus I^{\prime}|$ for some $(r^{\prime},I^{\prime})\in\mathcal{E}^{\mathrm{r}}\cup\{(0,\emptyset)\}$. First assume that $I\cap I^{\prime}\neq\emptyset$. Let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{\mathrm{r}}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{\mathrm{r}}$ contained in $I\cap I^{\prime}$. Then by (E3.2) $0+r^{\prime}\geq r_{3}+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|$. Then $r^{\prime}+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$. Hence $r([i,j-1])=r_{3}+|[i,j-1]\setminus I_{3}|=r([i,j])$, where the last equality holds since $j\in I_{3}$. Assume now that $I\cap I^{\prime}\neq\emptyset$. Similarly, let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{\mathrm{r}}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{\mathrm{r}}$ contained in the minimal cyclic interval containing $I\cup I^{\prime}$ in $[i,j]$ $C$, minus $I\cup I^{\prime}$. Then by (E3.1), $0+r^{\prime}\geq r_{3}-r_{4}-|C\setminus I_{4}|$. Then $r^{\prime}+|[i,j-1]\setminus I^{\prime}|\geq r_{3}-r_{4}-|C\setminus I_{4}|+|[i,j-1]\setminus I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$ and as before we get that $r([i,j])=r([i,j-1])$ and we get that $r$ and $\operatorname{rk}$ agree on $[i,j]$.
Case 2.2. Suppose $\pi(j)>j$. Since rk ([$\textrm{rk }([i,j-1])=\textrm{rk }([i,j])$, by Proposition 2.7, there is some $h\in[i+1,j-1]$ such that $\pi(h)=j$, i.e. $r([h,j])=r([h+1,j])=r([h,j-1])$. By (r4), there exists some $(r,I)$ such that $[h,j]\subseteq I$ and $r=r([h,j])=\textrm{rk }([h,j])$. Let $(r^{\prime},I^{\prime})\in\mathcal{E}^{\mathrm{r}}\cup\{(0,\emptyset)\}$ be such that $r([i,j-1])=r^{\prime}+|[i,j-1]\setminus I^{\prime}|$. Suppose $I\cap I^{\prime}\neq\emptyset$, the other case will be equivalent. Let $(r_{3},I_{3})$ be the minimal element in $\mathcal{E}^{\mathrm{r}}$ containing $I\cup I^{\prime}$ and let $(r_{4},I_{4})$ be the maximal element from $\mathcal{E}^{\mathrm{r}}$ contained in $I\cap I^{\prime}$. Then by (E3), $r^{\prime}+|[i,j-1]\setminus I^{\prime}|=r^{\prime}+|[i,j-1]\setminus(I\cup I^{\prime})|+|I\setminus I^{\prime}|\geq r^{\prime}+|[i,j-1]\setminus I_{3}|+|I\setminus I^{\prime}|\geq r+r^{\prime}-|I\cap I^{\prime}|+|[i,j-1]\setminus I_{3}|\geq r_{3}+|[i,j-1]\setminus I_{3}|+r_{4}+|(I\cap I^{\prime})\setminus I_{4}|-|I\cap I^{\prime}|\geq r_{3}+|[i,j-1]\setminus I_{3}|$. Hence $r([i,j-1])=r_{3}+|[i,j-1]\setminus I_{3}|=r_{3}+|[i,j]\setminus I_{3}|=r([i,j])$, since $j\in I_{3}$. ∎
@@
If $\mathcal{E}^{\mathrm{r}}$ is a family of labeled cyclic interval satisfying (E1), (E2), and (E3) we can recover the bounded affine permutation of the positroid defined by $\mathcal{E}^{\mathrm{r}}$ by applying Algorithm 1 to $E^{\mathrm{r}}=\{(r,(i,j))\mid(r,[i,j+i-1])\in\mathcal{E}^{\mathrm{r}}\}$.